Suppose Avon and Nova stocks have volatilities of 52% and 23%, respectively, and they are perfectly negatively correlated. What portfolio of these two stocks has zero risk?
Answer #1

For negatively correlated stocks,

Standard Deviation = 0.52(w1) + 0.23(w2)

Here,

0 = 0.52(w) + 0.23(1 - w)

w = -79.31%

So,

Weight of Avon Stock = -79.32%

Weight of Nova Stock = 179.32%

Negative sign denote one need to short sell the stock.
Answer #2

SDp = w1SD1 - w2SD2 = 0

w1 = SD2/(SD1+SD2)

w2 = SD1/(SD1+SD2)


@SDp=0, there is no such portfolio that exists in the world

source: class
answered by: Anatora Aurorua